OpenStudy (anonymous):

What is the equation of a line that passes through the point (6,3) and is perpendicular to a line with a slope -3/2

OpenStudy (anonymous):

so its y=2/3.....

hero (hero):

Hint: If the two lines are perpendicular to each other, the product of their slopes will be -1. In other words: \(m_1m_2 = -1\)

hero (hero):

But remember the equation of a line is y = mx + b. The line will include the slope, but not just the slope by itself.

OpenStudy (anonymous):

y=2/3-1

hero (hero):

Please show the work you did to find the equation of the line.

OpenStudy (anonymous):

well i found the reciprocal on the -3/2

OpenStudy (anonymous):

and then i guessed on the negative 1

hero (hero):

You cannot guess with this. You have to show the work you did to get your results. You have to show that you know what to do in order to get the correct result.